A company's production has a process range of 20 inches when operating properly. However, the company is concerned about whether the process is still in control and wants to develop a control chart to measure it. They collected 12 daily samples of 25 observations each with the following sample ranges: 18.5,19.2,20.4,22.8,26.9,27.5,24.5,23.4,21.2,15.5,17.7,25.5

What are the upper and lower control limits respectively for the R chart?

30.82;9.18

38.46;1.54

21.54;18.46

40.00;0.00

24.62;15.38
Answer #1

Total number of samples = 263.1

Therefore, R-bar = 263.1 ÷ 12

= 21.93

To calculate the upper and lower control limits of R chart -

UCL = R-bar × D4

= 21.93 × 1.72

= 37.72.

LCL = R-bar × D3

= 21.93 × 0.28

= 6.14.

Hope this helps :)